OpenStudy (halfdeafdarling):

Need help can someone explain? I know it is the inverse I need to figure out, but I am confused. Given f(x)= 3x-1/2, solve for f^-1(4)

OpenStudy (fibonaccichick666):

ok, so can you first tell me what the inverse function is?

OpenStudy (fibonaccichick666):

if not, do you know how to solve for it?

OpenStudy (halfdeafdarling):

the f\[f^-1(4) ?\]

OpenStudy (halfdeafdarling):

sorry little confused the way it is explain doesn't expand on this

OpenStudy (halfdeafdarling):

Because the f-1 (x) is usally that i was expecting to go back and do f(4)=3(4)-1/2

OpenStudy (halfdeafdarling):

so i think im getting myself more backwards than I should be

OpenStudy (halfdeafdarling):

usually* sorry if spelling is wonky hah

OpenStudy (fibonaccichick666):

ok, so in order to find the inverse function(and points), all we have to do is switch the x and y. So let's do the whole find the function first so you get to see that

OpenStudy (fibonaccichick666):

soo you have \[ f(x)= \frac{3x-1}{2}\] which is \[ y=\frac{ 3x-1}{2}\]

OpenStudy (halfdeafdarling):

oh yeah I know how to do that it would be instead of the f(x)=3x-1/2 it would be y=3x-1 but then have to switch it to x=3y-1 ?

OpenStudy (fibonaccichick666):

so, we switch it like you said, and we are here: \[ x= \frac{3y-1}{2}\]

OpenStudy (anonymous):

Switch x and y. Then solve for y. Insert 4 into the equation.

OpenStudy (fibonaccichick666):

now, can you solve for y?

OpenStudy (halfdeafdarling):

so then it would be 4=3y-1/2 going to * 2 so it would then be 8=3y-1 +1 being 9=3y/3 getting 3=y or y=3?

OpenStudy (fibonaccichick666):

uhm, well, you can do it that way. nice job. So you found the point (4,3) so, let's see what happens when we plug 3 into the original equation

OpenStudy (fibonaccichick666):

so, we'd have \(y=(3(3)-1)/2\)

OpenStudy (fibonaccichick666):

Do we get y=4?

OpenStudy (halfdeafdarling):

this is where I was getting confused when doing (3(3)-1)/2 i am getting 5/2

OpenStudy (fibonaccichick666):

what's 3 times 3?

OpenStudy (halfdeafdarling):

9....

OpenStudy (halfdeafdarling):

im dumb xD

OpenStudy (fibonaccichick666):

haha, you're not. Just learning something new. You got this

OpenStudy (halfdeafdarling):

oooohkay so thats where it was 4 --- i kept doing 3*2 thinking it was like that --- so 8/2 would be 4

OpenStudy (jdoe0001):

If you shut your door to all errors truth will be shut out. ~~ Rabindranath Tagore (1861-1941) ~~

OpenStudy (fibonaccichick666):

yea, so now, we have checked our answer. If it didn't plug back in and give us 4, we would have known that we did something wrong.

OpenStudy (halfdeafdarling):

okay so final answer to all this is that 3, and doing the original equation and plugging it in will allow us to check our equation? Thank you, it cleared it up

OpenStudy (fibonaccichick666):

yep yep. np :)

OpenStudy (fibonaccichick666):

May I ask you something?

OpenStudy (halfdeafdarling):

ask away!

OpenStudy (fibonaccichick666):

Are you Deaf?

OpenStudy (fibonaccichick666):

Do you use ASL?

OpenStudy (halfdeafdarling):

I am half deaf on my right side and mechanically deaf on my left Surgery caused me to be like this

OpenStudy (halfdeafdarling):

and I am wanting to get into ASL but its hard to do studying or get training when having to go to medical centers and trying to balance a full schedule of schooling and musicianship

OpenStudy (fibonaccichick666):

Ah, ok, I was wondering. I would like to get my masters or doctorate in Deaf education(I'm fluent in ASL), so I find it helpful to know what writing style is useful for students to understand(since most have english as their second language)

OpenStudy (fibonaccichick666):

A lot of schools offer ASL as a language credit now. You might be able to take it to complete your degree.

OpenStudy (halfdeafdarling):

I think writing out things have always been easier for me to due cause I also have movement problems in m hands but typing helps because i can delete anything. I twitch and when trying to sign my signing can be misread. Ever since i was little i would always write something down if I was worried about not hearing, or not being able to be undersstood

OpenStudy (halfdeafdarling):

I do full online as of right now, I was going to take it for my next semester course but things caused a schedule change and as of right now my aunt (who is also fluent in ASL and teaches) on free days trains me

OpenStudy (fibonaccichick666):

That's a good method for when you want to ensure your message is relayed. And that's awesome. I love hearing about people who have family who sign too. So many of the students I observed did not have that, so they dreaded going home on weekends.

OpenStudy (halfdeafdarling):

Oh by far, a friend of mine who was born completely deaf, would sit and start drawing which was always something we both loved, always were scared if we go to our families and can't communicate with them, they will get more frustrated and not want to help or cause household problems. Thankfully before I had left the school her family and her enrolled in a family course for ASL with students that doesn't have the family members, along with the school providing ASL workers around the school to always help teachers or other students understand us.

OpenStudy (halfdeafdarling):

I am also happy to hear you (pun hah) are going through the course for ASL, I wish you all of the luck!

OpenStudy (fibonaccichick666):

That's good at least. And thanks, it will be a couple of years, but I'm hoping to go to Gallaudet for it. I'd love to teach there. Thanks for answering my question!

OpenStudy (halfdeafdarling):

And thank you for answering mine! Hope to talk again you seem to be a very nice human being!

OpenStudy (fibonaccichick666):

You too! Feel free to tag me any time!